# Probing Reheating Phase via Non-Helical Magnetogenesis and Secondary Gravitational Waves

### v1

@@ III Generation of Gravitational Waves:
where ‘$\eta$’ is the conformal time and $h_{ij}$ is the traceless tensor, i.e., $\partial^{i}h_{ij}=h^{i}_{i}=0$. In Fourier space, the equation of motion for the gravitational wave amplitude ‘$h$’, for either polarization $h^{+}$ or $h^{-}$, becomes Caprini and Durrer (2001); Sorbo (2011); Caprini and Sorbo (2014); Ito and Soda (2017); Sharma et al. (2020); Okano and Fujita (2021):
@@
[…]
@@ III.1 Generation and Evolution of Primary Gravitational Waves
After inflation, the universe undergoes different phases, and the initial fluctuations are modified as they pass through these phases, encoding information about the universe’s history. We consider a non-instantaneous reheating scenario, where the scale factor evolves as $a(\eta)\simeq a_{\mathrm{e}}\left(\eta/\eta_{\rm e}\right)^{\delta/2}$, with $\delta=4/(1+3w_{re})$, where $w_{re}$ is the average equation of state (EoS) during reheating. Solving the homogeneous part of Eq. (33) during the reheating phase, the tensor fluctuations $h_{\textbf{k}}$ in the super-horizon approximation $(k<k_{\rm e})$ are given by
@@
$\mathcal{P}(k,\eta>\eta_{\rm re})\simeq\frac{1}{2(k\eta)^{2}}\left(1+\left(\frac{k}{k_{\mathrm{re}}}\right)^{2}\right)\mathcal{P}(k,\eta_{\rm re}).$ (44)
Here, $\mathcal{P}(k,\eta_{\rm re})$ represents the total tensor power spectrum at the end of reheating. For modes that remain outside the horizon at the end of reheating $(k<k_{\mathrm{re}})$, the tensor power spectrum in the radiation-dominated era scales proportionally to the inflationary tensor power spectrum, i.e., $\mathcal{P}(k,\eta>\eta_{\rm re})\propto\mathcal{P}(k,\eta_{\rm e})$. However, for modes that re-entered the horizon before reheating ended $(k>k_{\mathrm{re}})$, the tensor power spectrum is affected by the reheating history, scaling as $\mathcal{P}(k>k_{\mathrm{re}})\propto x_{re}^{2l+1}\mathcal{P}(k,\eta_{\rm e})$, where $x_{re}=k\eta_{\rm re}$.
@@ III.2 Generation of Secondary Gravitational Waves
Electromagnetic (EM) fields in the early universe offer a compelling secondary source of gravitational waves (GWs), capable of amplifying the GW spectrum beyond inflationary contributions. There are three different phases where we can expect that the Electromagnetic field can have the potential to generate the GW signal enough that it can be passed through different future observatories. We consider that electromagnetic fields are generated during inflation (discussed in sec.(II)), which has significant strength (depending on the model parameters). The electromagnetic fields are one of the key sources that can generate gravitational waves during inflation as well as the post-inflationary era. Now, to see the contributions due to the EM field during inflation, we are going to use the Green’s function method like Maiti et al. (2024). During inflation in a de Sitter background, the Green’s function is well established Cook and Sorbo (2012),
@@
$\mathcal{P}_{T,s}^{\mathrm{inf}}(k,\eta_{\rm e})=2\left(\frac{H_{{}_{\mathrm{I}}}}{M_{{}_{\mathrm{P}}}}\right)^{4}\mathcal{C}_{{}_{\rm B/E}}^{2}(n_{{}_{\mathrm{B}/\mathrm{E}}})\left(\frac{k}{k_{\rm e}}\right)^{2n_{{}_{\mathrm{B}/\mathrm{E}}}}\mathcal{I}_{\rm inf}^{2}~{}\mathcal{F}_{n_{{}_{\mathrm{B}/\mathrm{E}}}}(k),$ (47)
where $\mathcal{I}_{\rm inf}=(1-(k_{\rm e}/k_{\rm min})^{n_{{}_{\mathrm{B}/\mathrm{E}}}})/3n_{{}_{\mathrm{B}/\mathrm{E}}}$ and we also define,
$\mathcal{C}_{{}_{\rm B/E}}(n_{{}_{\mathrm{B}/\mathrm{E}}})=\frac{2^{5-n_{{}_{\mathrm{B}/\mathrm{E}}}}~{}\Gamma^{2}\left(\frac{5-n_{{}_{\mathrm{B}/\mathrm{E}}}}{2}\right)}{8\pi^{3}}$ (48)
@@
$\mathcal{F}_{n_{{}_{\mathrm{B}/\mathrm{E}}}}(k)=\frac{8}{3n_{{}_{\mathrm{B}/\mathrm{E}}}}\left(1-\left(\frac{k_{\rm min}}{k}\right)^{n_{{}_{\mathrm{B}/\mathrm{E}}}}\right)+\frac{56}{15(2n_{{}_{\mathrm{B}/\mathrm{E}}}-3)}\left(\left(\frac{k_{\mathrm{e}}}{k}\right)^{2n_{{}_{\mathrm{B}/\mathrm{E}}}-3}-1\right).$ (49)
Here $k_{\rm min}$ is the lowest possible scale. As in this case we consider a non-helical magnetogenesis model discussed in sec.(II), depending on the nature of the coupling, one of the fields, either the magnetic field or the electric field, is the dominating component. So, if we change the nature of the coupling, the final spectrum behavior will also change. But at the end of inflation, it is very hard to distinguish, as for this type of coupling we have the same behavior when we change the sign of the coupling constant $n$ (see in Eqs. (16)(17)). However, when the universe evolves through different phases, there will be a significant change in the overall GW spectrum at present, which will be discussed in detail later. But at the end of inflation, we have found that at large scales, such as those associated with the CMB, the primary production is more dominant compared to the secondary as $\mathcal{P}_{\rm v}^{\rm inf}(k)\propto(H_{{}_{\mathrm{I}}}/M_{\rm P})^{2}$, whereas $\mathcal{P}^{\rm inf}_{\rm s}(k)\propto(H_{{}_{\mathrm{I}}}/M_{\rm P})^{4}$, where $\mathcal{P}_{\rm v}^{\rm inf}(k)$ is the contribution due to quantum fluctuations and $\mathcal{P}^{\rm inf}_{\rm s}(k)$ is the contribution due to EM fields.
After inflation, we have the reheating phase, where all the fundamental particles are produced. Depending on the conductivity during this time, the nature of the GWs produced due to EM fields also changes. Here, we consider two extreme scenarios: either we consider that during the reheating phase the effective conductivity of our universe is very low (for simplicity, we consider zero conductivity), or we can consider that initial production of charge particles were enough that during this phase the conductivity was very high (for simplicity, we consider infinite conductivity scenario). When we consider zero electric conductivity, both the electric and magnetic fields contribute, whereas in the other case, electric fields vanish, and only the magnetic field contributes to secondary gravitational waves. During the reheating phase, the Green’s function can be written as  Maiti et al. (2024),
@@
In general, we can define the tensor power spectrum generated during reheating due to the EM field, as defined at the end of reheating, by the following expression
$\mathcal{P}_{T,s}^{\mathrm{re}}=\frac{2H_{{}_{\mathrm{I}}}^{4}}{M_{\rm P}^{4}}\left(\frac{k}{k_{\mathrm{e}}}\right)^{2(\delta-2)}\mathcal{C}_{{}_{\rm B/E}}^{2}(n_{{}_{\mathrm{B}/\mathrm{E}}})\left(\frac{k}{k_{\rm e}}\right)^{2n_{{}_{\mathrm{B}/\mathrm{E}}}}\mathcal{C}_{\mathrm{m,re}}^{2}(x_{re},x_{\mathrm{end}})\mathcal{F}_{n_{{}_{\mathrm{B}/\mathrm{E}}}}(k),$ (51)
@@
[…]
If we analyze the spectrum, we find that due to the non-instantaneous reheating phase, the overall spectrum has a term like $(k_{\mathrm{re}}/k_{\rm e})^{2(\delta-2)}$. We recall that $\delta={4}/{(1+3w_{re})}$. So, depending on the nature of the reheating scenarios, either the overall spectrum gets enhanced (for $w_{re}>1/3$) or suppressed (for $w_{re}<1/3$). Although we have also found that for those modes which are outside the horizon $k<k_{\mathrm{re}}$, the secondary spectrum always goes as $\mathcal{P}^{\rm re}_{\rm s}(k<<k_{\mathrm{re}},\eta_{\rm re})\propto k^{2n_{{}_{\mathrm{B}/\mathrm{E}}}}$, whereas for those modes which are inside the horizon, it goes as $\mathcal{P}^{\rm re}_{\rm s}(k>>k_{\mathrm{re}},\eta_{\rm re})\propto k^{2n_{{}_{\mathrm{B}/\mathrm{E}}}-|n_{w}|-2}$. Here, $n_{w}=2(1-3w_{re})/(1+3w_{re})$.
During the radiation-dominated era, the electrical conductivity of our universe was very high, and the only existing component was the magnetic field, which produces GWs. It is well known that such production is only effective up to the neutrino decoupling era. After neutrinos decouple from the thermal bath, they effectively cancel the anisotropy due to magnetic fields Lewis (2004).
Thus, the tensor power spectrum due to the magnetic field after reheating, defined at the time of neutrino decoupling, is governed by the expression Maiti et al. (2024)
$\mathcal{P}^{\lambda}_{{}_{\rm ra,SEC}}(k,\eta_{\nu})=\frac{2H_{{}_{\mathrm{I}}}^{4}}{k^{2}\eta^{2}_{\nu}M_{\rm P}^{4}}\left(\frac{k_{\mathrm{re}}}{k_{\rm e}}\right)^{2(\delta-2)}\mathcal{C}_{{}_{\rm B/E}}^{2}(n_{\mathrm{B}})\left(\frac{k}{k_{\rm e}}\right)^{2(4-|2n|)}{\mathcal{I}_{\rm ra}}^{2}(x_{\nu},x_{re})\mathcal{F}_{n_{{}_{\rm B}}}(k)$ (53)
Here, $\mathcal{I}_{\rm ra}(x_{\nu},x_{re})$ is defined as Maiti et al. (2024)
[…]
@@
Here, we have seen that there is an overall enhancement factor associated with the tensor power spectrum produced during this phase. This extra enhancement factor comes from the phase of reheating, as during reheating the background energy density dilutes as $a^{-3(1+w_{re})}$, whereas the magnetic field always decays as $a^{-4}$. Depending on the EoS, where we have $w_{re}<1/3$ or $w_{re}>1/3$, the contribution can either be suppressed or enhanced. In this specific kind of magnetogenesis model, we get the effective contribution in the frequency range $f_{\nu}<f<f_{\rm re}$ only if we consider $w_{re}>1/3$ with $n>1/2$. We recall that here, $n$ is the coupling constant associated with the magnetogenesis model that we discussed in Sec.(II).
@@ III.3.2 Spectral Analysis of Secondary Gravitational Waves (SGWs)
In analyzing the spectrum of Secondary Gravitational Waves (SGWs) across a wide frequency range, we focus on contributions from magnetogenesis and ignore the primary inflationary GWs. Depending on the coupling parameter in the magnetogenesis model, either strong magnetic fields (for $n>1/2$) or strong electric fields (for $n<-1/2$) dominate. Additionally, the final spectral characteristics vary with the level of electrical conductivity, leading us to discuss the scenarios separately.
@@
Assuming zero electrical conductivity, both electric and magnetic fields are present during reheating. Depending on the coupling nature, SGWs are sourced primarily by electric fields (for $n<-1/2$) or magnetic fields (for $n>1/2$). In reheating scenarios where $w_{re}<1/3$, the spectral behavior is similar for both cases (either $n>1/2$ or $n<-1/2$), differing only in whether the contribution originates from the magnetic or electric field. The SGW spectrum in this regime is given by Maiti et al. (2024)
@@
$\Omega_{\mathrm{gw}}^{\mathrm{sec}}(k)h^{2}\simeq$  $\,2.26\times 10^{-26}\left(\frac{\Omega_{R}h^{2}}{4.3\times 10^{-5}}\right)\left(\frac{H_{{}_{\mathrm{I}}}}{10^{-5}M_{\rm P}}\right)^{4}\left(\frac{k_{\mathrm{re}}}{k_{\rm e}}\right)^{2n_{{}_{\mathrm{B}/\mathrm{E}}}}\mathcal{I}_{\rm inf}^{2}\mathcal{C}_{\beta}^{2}(n_{{}_{\mathrm{B}/\mathrm{E}}})\,\mathcal{F}_{\beta}(k)$
×$\times\begin{cases}\left(\frac{k}{k_{\mathrm{re}}}\right)^{2n_{{}_{\mathrm{B}/\mathrm{E}}}},&k_{*}<k<k_{\mathrm{re}},\\ \left(\frac{k}{k_{\mathrm{re}}}\right)^{2n_{{}_{\mathrm{B}/\mathrm{E}}}-|n_{w}|},&k_{\mathrm{re}}<k<k_{\rm e}.\end{cases}$ (58)
For reheating scenarios with $w_{re}>1/3$, on the other hand, the final spectrum changes depending on whether $n>1/2$ or $n<-1/2$. When $n>1/2$, the dominant component is the magnetic field, yielding the spectrumMaiti et al. (2024)
@@
$\Omega_{\rm gw,pinf}^{\rm sec}(k)h^{2}\simeq$  $\,2.26\times 10^{-26}\left(\frac{\Omega_{R}h^{2}}{4.3\times 10^{-5}}\right)\left(\frac{H_{{}_{\mathrm{I}}}}{10^{-5}M_{\rm P}}\right)^{4}\left(\frac{k_{\mathrm{re}}}{k_{\rm e}}\right)^{2(n_{w}+n_{{}_{\mathrm{B}/\mathrm{E}}})}\mathcal{C}_{\beta}^{2}(n_{\mathrm{B}})\,\mathcal{F}_{\beta}(k)$
@@
×$\times\begin{cases}\mathcal{A}_{1}\left(\frac{k}{k_{\mathrm{re}}}\right)^{2n_{\mathrm{B}}},&k_{*}<k<k_{\rm ssb},\\ \mathcal{I}_{2}(k)\left(\frac{k_{\mathrm{re}}}{k_{\nu}}\right)^{2}\left(\frac{k}{k_{\mathrm{re}}}\right)^{2n_{\mathrm{B}}+2},&k_{\rm ssb}<k<k_{\nu},\\ \mathcal{I}_{2}(k)\left(\frac{k}{k_{\mathrm{re}}}\right)^{2n_{\mathrm{B}}},&k_{\nu}\leq k\leq k_{\mathrm{re}},\\ \mathcal{A}_{2}\left(\frac{k}{k_{\mathrm{re}}}\right)^{2n_{\mathrm{B}}-|n_{w}|},&k>k_{\mathrm{re}}.\end{cases}$ (59)
Here, $k_{\nu}$ denotes the wavenumber that re-enters the horizon at the time when neutrinos decouple from the thermal bath. The wavenumber $k_{\rm ssb}$ corresponds to the scale at which production during the radiation-dominated era surpasses that during the reheating phase. It can be approximately defined as $k_{\rm ssb}=\mathcal{A}_{1}^{1/2}{k_{\nu}}/{\ln(k_{\mathrm{re}}/k_{\nu})}\simeq 10^{-1}k_{\nu}.$ Conversely, for couplings with $n<-1/2$, the electric field is dominant, and after reheating, only a weak residual magnetic field remains. Consequently, contributions from the magnetic field are negligible in comparison to the primary electric field contributions. The SGW spectrum in this case primarily follows
$\Omega_{\rm gw,pinf}^{\rm sec}(k)h^{2}\simeq$  $\,2.26\times 10^{-26}\left(\frac{\Omega_{R}h^{2}}{4.3\times 10^{-5}}\right)\left(\frac{H_{{}_{\mathrm{I}}}}{10^{-5}M_{\rm P}}\right)^{4}\left(\frac{k_{\mathrm{re}}}{k_{\rm e}}\right)^{2(n_{w}+n_{\mathrm{E}})}\mathcal{C}_{\beta}^{2}(n_{\mathrm{E}})\,\mathcal{F}_{n_{\mathrm{E}}}(k)$
×$\times\begin{cases}\mathcal{A}_{1}\left(\frac{k}{k_{\mathrm{re}}}\right)^{2n_{\mathrm{E}}},&k_{*}<k<k_{\mathrm{re}},\\ \mathcal{A}_{2}\left(\frac{k}{k_{\mathrm{re}}}\right)^{2n_{\mathrm{E}}-|n_{w}|},&k>k_{\mathrm{re}}.\end{cases}$ (60)
For $n>1/2$, the overall spectrum includes additional spectral breaks, as magnetic fields contribute to GW production up to the neutrino decoupling era. In contrast, for $n<-1/2$, the high-conductivity post-reheating environment washes out the electric field, leaving only a small magnetic field contribution, making the SGW signal significantly weaker than primary contributions. Note that the quantities $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are listed in Appendix A.

### v2

@@ III Generation of Gravitational Waves:
where $\eta$’ is the conformal time and $h_{ij}$ is the traceless tensor, i.e., $\partial^{i}h_{ij}=h^{i}_{i}=0$. In Fourier space, the equation of motion for the gravitational wave amplitude ‘$h$’, for either polarization $h^{+}$ or $h^{-}$, becomes Caprini and Durrer (2001); Sorbo (2011); Caprini and Sorbo (2014); Ito and Soda (2017); Sharma et al. (2020); Okano and Fujita (2021):
@@
$\mathcal{P}^{\lambda}_{{}_{\rm SEC}}(k,\eta)$ $=\frac{2}{M_{\rm P}^{4}}\left[\int_{\eta_{i}}^{\eta}d\eta_{1}a^{2}(\eta_{1})\mathcal{G}_{k}(\eta_{f},\eta_{1})\mathcal{P}_{\rm B/E}(k,\eta_{1})\right]^{2}\times\int_{0}^{\infty}\frac{dq}{q}\int_{-1}^{1}d\mu\frac{(q/k)^{n_{{}_{\mathrm{B}/\mathrm{E}}}}f(\mu,\beta,\lambda)}{\left[1+\left(\frac{q}{k}\right)^{2}-2\mu\frac{q}{k}\right]^{\frac{3-n_{{}_{\mathrm{B}/\mathrm{E}}}}{2}}}$ (40)
@@ III.1 Generation and Evolution of Primary Gravitational Waves
After inflation, the universe undergoes different phases, and the initial fluctuations are modified as they pass through these phases, encoding information about the universe’s history. We consider a non-instantaneous reheating scenario, where the scale factor evolves as $a(\eta)\simeq a_{\rm end}\left(\eta/\eta_{\rm e}\right)^{\delta/2}$, with $\delta=4/(1+3w_{re})$, where $w_{re}$ is the average equation of state (EoS) during reheating. Solving the homogeneous part of Eq. (35) during the reheating phase, the tensor fluctuations $h_{\textbf{k}}$ in the super-horizon approximation $(k<k_{\rm e})$ are given by
@@
$\mathcal{P}_{\rm T}(k,\eta>\eta_{\rm re})\simeq\frac{1}{2(k\eta)^{2}}\left(1+\left(\frac{k}{k_{\mathrm{re}}}\right)^{2}\right)\mathcal{P}_{\rm T}(k,\eta_{\rm re}).$ (46)
Here, $\mathcal{P}_{\rm T}(k,\eta_{\rm re})$ represents the total tensor power spectrum at the end of reheating. For modes that remain outside the horizon at the end of reheating $(k<k_{\mathrm{re}})$, the tensor power spectrum in the radiation-dominated era scales proportionally to the inflationary tensor power spectrum, i.e., $\mathcal{P}_{\rm T}(k,\eta>\eta_{\rm re})\propto\mathcal{P}_{\rm T}(k,\eta_{\rm e})$. However, for modes that re-entered the horizon before reheating ended $(k>k_{\mathrm{re}})$, the tensor power spectrum is affected by the reheating history, scaling as $\mathcal{P}_{\rm T}(k>k_{\mathrm{re}})\propto x_{re}^{2l+1}\mathcal{P}_{\rm T}(k,\eta_{\rm e})$, where $x_{re}=k\eta_{\rm re}$.
@@ III.2 Generation of Secondary Gravitational Waves
Electromagnetic (EM) fields in the early Universe provide a compelling secondary source of gravitational waves (GWs), with the potential to enhance the GW spectrum beyond the standard inflationary background. There are three distinct cosmological phases during which EM fields can source GWs: during inflation, during reheating, and in the subsequent radiation-dominated era. In this work, we focus on scenarios where EM fields are generated during inflation (as discussed in Sec. II), potentially attaining significant amplitudes depending on the model parameters.
In such models, EM fields serve as key sources for GWs throughout both the inflationary and post-inflationary epochs. However, successful magnetogenesis in these scenarios—avoiding strong coupling and backreaction problems—requires the coupling parameter to lie within the range 2.2<n<0-2.2<n<0. Within this regime, the electric field typically dominates on large scales, implying that the corresponding tensor perturbations sourced by the EM field are predominantly generated by the electric component.
Moreover, the post-inflationary evolution plays a crucial role in determining the present-day magnetic field strength. Specifically, in the case of vanishing electrical conductivity during reheating, the induction effect leads to significant amplification of large-scale magnetic fields, even for $n<0$ couplings, while still evading strong coupling and backreaction issues. Therefore, in this section, we focus on scenarios that produce large-scale magnetic fields within the allowed parameter range and remain free from theoretical inconsistencies.
To compute the tensor power spectrum sourced by the electric field during inflation, we solve Eq. (35) using the Green’s function method, following the approach of Maiti et al. (2024). In a de Sitter background, the Green’s function is well established and has been extensively studied in the literature (see, for example, Cook and Sorbo (2012)).
@@
$\mathcal{P}_{T,s}^{\mathrm{inf}}(k,\eta_{\rm e})=2\left(\frac{H_{{}_{\mathrm{I}}}}{M_{{}_{\mathrm{P}}}}\right)^{4}\mathcal{C}_{\rm E}^{2}(n_{\mathrm{E}})\left(\frac{k}{k_{\rm end}}\right)^{2n_{\mathrm{E}}}\mathcal{I}_{\rm inf}^{2}~\mathcal{F}_{n_{\rm E}}(k),$ (49)
where $\mathcal{I}_{\rm inf}\simeq 1/3n_{\mathrm{E}}$ and we also define,
$\mathcal{C}_{\rm E}(n_{\mathrm{E}})=\frac{2^{5-n_{\mathrm{E}}}~\Gamma^{2}\left(\frac{5-n_{\mathrm{E}}}{2}\right)}{8\pi^{3}}$ (50)
@@
$\mathcal{F}_{n_{\rm E}}(k)=\frac{8}{3n_{\mathrm{E}}}\left(1-\left(\frac{k_{*}}{k}\right)^{n_{\mathrm{E}}}\right)+\frac{56}{15(2n_{\mathrm{E}}-3)}\left(\left(\frac{k_{\rm e}}{k}\right)^{2n_{\mathrm{E}}-3}-1\right).$ (51)
In this scenario, since the dominant contribution arises entirely from the electric field, the tensor power spectrum sourced by the electric field is expected to scale as $\mathcal{P}_{T,s}^{\mathrm{inf}}(k)\propto\left({H_{{}_{\mathrm{I}}}}/{M_{\rm P}}\right)^{4}\left({k}/{k_{\rm e}}\right)^{n_{\mathrm{E}}}$, given that the electric field power spectrum behaves as $\mathcal{P}_{E}(k)\propto H_{{}_{\mathrm{I}}}^{4}\left({k}/{k_{\rm e}}\right)^{n_{\mathrm{E}}}$. In contrast, the tensor power spectrum originating from the vacuum quantum fluctuations of spacetime follows $\mathcal{P}_{T,v}^{\mathrm{inf}}(k)\propto\left({H_{{}_{\mathrm{I}}}}/{M_{\rm P}}\right)^{2}$.
Typically, for blue-tilted electric spectra with $n_{\mathrm{E}}>0$, the vacuum contribution dominates on large (CMB) scales, since the electric field power is concentrated at small scales. However, in the case of $n_{\mathrm{E}}<0$ (corresponding to n<2n<-2), the electric field becomes more red-tilted, concentrating its energy on large scales. This opens the possibility that, at CMB scales, the secondary tensor modes sourced by the electric field could exceed the primary vacuum contribution.
After inflation, all fundamental particles are produced through the decay of the inflaton field. The nature of gravitational wave (GW) production in the subsequent reheating era depends sensitively on the electrical conductivity of the Universe. As previously discussed, in the limit of negligible conductivity, both electric and magnetic fields are present and evolve freely. In contrast, in the case of large or infinite conductivity, the electric field is quickly suppressed due to the high response of charged particles, leaving only the magnetic field. Additionally, in the zero-conductivity limit, energy transfer from the electric to the magnetic field via the Faraday induction mechanism significantly enhances the magnetic field amplitude. However, such enhancement does not occur in the high-conductivity regime, resulting in much weaker magnetic fields.
Consequently, in the infinite conductivity limit, the secondary GW production sourced by the magnetic field is expected to be suppressed, particularly at large scales. Since we are interested in scenarios that lead to successful magnetogenesis without encountering strong coupling or backreaction issues, we focus on the case where conductivity during reheating is negligible and the coupling parameter lies within the range 2.2<n<0-2.2<n<0. In this regime, the dominant post-inflationary contribution to the stochastic gravitational wave background (SGWB) arises from the electric field.
During the reheating phase, the corresponding Green’s function can be expressed as Maiti et al. (2024)
@@
Since the tensor power spectrum is sourced by the electric field, the tensor spectrum generated during reheating due to the electromagnetic field, evaluated at the end of reheating, can be generally expressed as
$\mathcal{P}_{T,s}^{\mathrm{re}}=\frac{2H_{{}_{\mathrm{I}}}^{4}}{M_{\rm P}^{4}}\left(\frac{k}{k_{\rm e}}\right)^{2(\delta-2)}\mathcal{C}_{\rm E}^{2}(n_{\mathrm{E}})\left(\frac{k}{k_{\rm end}}\right)^{2n_{\mathrm{E}}}\mathcal{C}_{\mathrm{m,re}}^{2}(x_{re},x_{\mathrm{end}})\mathcal{F}_{n_{\rm E}}(k),$ (53)
@@
$\mathcal{C}_{\mathrm{m,re}}(x_{re},x_{\mathrm{end}})=\int_{x_{\mathrm{end}}}^{x_{re}}dx_{1}\,x_{1}^{-\delta}\mathcal{G}_{\textbf{k}}(x_{re},x_{1}),$ (54)
Upon analyzing the spectrum, we find that the presence of a non-instantaneous reheating phase introduces an overall factor of the form $\left({k_{\mathrm{re}}}/{\kappa_{\rm end}}\right)^{2(\delta-2)}$, where $\delta={4}/{(1+3w_{re})}$. This indicates that the amplitude of the spectrum can either be enhanced or suppressed depending on the equation of state during reheating: it is enhanced for $w_{re}>1/3$ and suppressed for $w_{re}<1/3$.
Moreover, we observe that for modes that remain outside the horizon during reheating ($k<k_{\mathrm{re}}$), the secondary tensor spectrum behaves as $\mathcal{P}^{\rm re}_{\rm s}(k\ll k_{\mathrm{re}},\eta_{\rm re})\propto k^{2n_{\mathrm{E}}}$. In contrast, for modes that re-enter the horizon during reheating ($k\gg k_{\mathrm{re}}$), the spectrum scales as $\mathcal{P}^{\rm re}_{\rm s}(k\gg k_{\mathrm{re}},\eta_{\rm re})\propto k^{2n_{\mathrm{E}}-|n_{w}|-2},$ where $n_{w}={2(1-3w_{re})}/{(1+3w_{re})}$.
For this specific choice of coupling, there are no significant additional contributions to the gravitational wave spectrum from the radiation-dominated era. This is because, following reheating, it is generally assumed that all fundamental particles are produced and the Universe becomes thermalized, resulting in a highly conductive plasma. As a consequence, the electric field component is rapidly suppressed. However, the magnetic field survives and evolves adiabatically, with its energy density decreasing solely due to the expansion of the background.
Therefore, the tensor power spectrum sourced by the magnetic field after reheating, evaluated at the time of neutrino decoupling, is governed by the expression Maiti et al. (2024).
$\mathcal{P}^{\lambda}_{{}_{\rm ra,SEC}}(k,\eta_{\nu})=\frac{2H_{{}_{\mathrm{I}}}^{4}}{k^{2}\eta^{2}_{\nu}M_{\rm P}^{4}}\left(\frac{k_{\mathrm{re}}}{k_{\rm end}}\right)^{2(\delta-2)}\mathcal{C}_{\rm B}^{2}(n_{\mathrm{B}})\left(\frac{k}{k_{\rm end}}\right)^{2n_{\mathrm{B}}}{\mathcal{I}_{\rm ra}}^{2}(x_{\nu},x_{re})\mathcal{F}_{n_{{}_{\rm B}}}(k)$ (55)
We recall that $n_{\mathrm{B}}=n_{\mathrm{E}}+2$ is the magnetic spectral index for $n<0$ magnetogenesis scenarios. Here, $\mathcal{I}_{\rm ra}(x_{\nu},x_{re})$ is defined as Maiti et al. (2024)
${\mathcal{I}_{\rm ra}(x_{\nu},x_{re})=\int_{x_{re}}^{x_{\nu}}dx_{1}\frac{\sin(x_{1}-x_{\nu})}{x_{1}}}.$ (56)
@@
In this context, we observe that the tensor power spectrum generated during this phase includes an overall enhancement factor. This enhancement originates from the dynamics of the reheating phase: while the background energy density dilutes as $a^{-3(1+w_{re})}$, the magnetic field energy density redshifts as $a^{-4}$. As a result, depending on the equation of state (EoS), the gravitational wave contribution from magnetic fields can either be enhanced (for $w_{re}>1/3$) or suppressed (for $w_{re}<1/3$).
Specifically, within the framework of this magnetogenesis model, a significant contribution to the tensor spectrum in the frequency range $f_{\nu}<f<f_{\rm re}$ arises only when $w_{re}>1/3$ and the coupling parameter satisfies $n>1/2$. Here, $n$ refers to the coupling constant associated with the magnetogenesis scenario discussed in Sec. II.
@@ III.3.2 Spectral Analysis of Secondary Gravitational Waves (SGWs)
In analyzing the spectrum of secondary gravitational waves (SGWs) over a broad frequency range, we concentrate on contributions arising from inflationary magnetogenesis, while neglecting the primary gravitational waves sourced by vacuum fluctuations. Two limiting scenarios are considered based on the conductivity of the Universe during reheating. In the first case, we assume zero electrical conductivity, where both electric and magnetic fields are present throughout the reheating phase. In the second scenario, we consider infinite electrical conductivity, in which only the magnetic field survives and evolves adiabatically after inflation.
@@
Under the assumption of negligible electrical conductivity during reheating, both electric and magnetic fields persist during this epoch. In the parameter range that avoids both strong coupling and backreaction issues, namely $0.4<n_{\mathrm{E}}<4$, the electric field typically dominates. Consequently, the SGW spectrum is predominantly sourced by the electric field during inflation and reheating. Although the magnetic field can contribute to SGW production in the subsequent radiation-dominated era (up to neutrino decoupling), the contribution from the reheating phase is generally dominant.
The shape of the resulting GW spectrum is sensitive to the background evolution during reheating. Specifically, for reheating scenarios with an equation of state parameter $w_{re}<1/3$, the present-day SGW spectrum sourced by the electromagnetic field is
@@
$\Omega_{\mathrm{gw}}^{\mathrm{sec}}(k)h^{2}\simeq$  $\,2.26\times 10^{-26}\left(\frac{\Omega_{R}h^{2}}{4.3\times 10^{-5}}\right)\left(\frac{H_{{}_{\mathrm{I}}}}{10^{-5}M_{\rm P}}\right)^{4}\left(\frac{k_{\mathrm{re}}}{k_{\rm end}}\right)^{2n_{\mathrm{E}}}\mathcal{I}_{\rm inf}^{2}\mathcal{C}_{\beta}^{2}(n_{\mathrm{E}})\,\mathcal{F}_{\beta}(k)$
$\times\begin{cases}\left(\frac{k}{k_{\mathrm{re}}}\right)^{2n_{\mathrm{E}}},&k_{*}<k<k_{\mathrm{re}},\\ \left(\frac{k}{k_{\mathrm{re}}}\right)^{2n_{\mathrm{E}}-|n_{w}|},&k_{\mathrm{re}}<k<k_{\rm e}.\end{cases}$ (60)
whereas for the reheating scenarios with $w_{re}>1/3$, the spectral behaviour due to EM field is
@@
$\Omega_{\rm gw,pinf}^{\rm sec}(k)h^{2}\simeq$  $\,2.26\times 10^{-26}\left(\frac{\Omega_{R}h^{2}}{4.3\times 10^{-5}}\right)\left(\frac{H_{{}_{\mathrm{I}}}}{10^{-5}M_{\rm P}}\right)^{4}\left(\frac{k_{\mathrm{re}}}{k_{\rm end}}\right)^{2(n_{w}+n_{\mathrm{E}})}\mathcal{C}_{\beta}^{2}(n_{\mathrm{E}})\,\mathcal{F}_{n_{\mathrm{E}}}(k)$
@@
$\times\begin{cases}\mathcal{A}_{1}\left(\frac{k}{k_{\mathrm{re}}}\right)^{2n_{\mathrm{E}}},&k_{*}<k<k_{\mathrm{re}},\\ \mathcal{A}_{2}\left(\frac{k}{k_{\mathrm{re}}}\right)^{2n_{\mathrm{E}}-|n_{w}|},&k>k_{\mathrm{re}}\end{cases}$ (61)
Here, we clearly observe an additional enhancement factor associated with reheating scenarios where $w_{re}>1/3$, which arises from the specific evolution of the background dynamics during the reheating phase. This enhancement originates from the differential dilution rates between the background energy density and the electromagnetic (EM) field energy density. For $w_{re}>1/3$, the background energy density redshifts more rapidly than the EM field, leading to a relative amplification of the SGW signal. The coefficients $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, which characterize this enhancement, are provided in Appendix A.